solve the following quadratic equaton for x
(2x+1)(x+3)=0

Answer by ewatrrr(24785)

 
Hi,
If (2x+1)(x+3)=0 ⇒
(2x+1)= 0
or
(x+3)=0
x = -1/2 or x = -3